OpenStudy (anonymous):

8x + 15 = 10x solve for x

OpenStudy (anonymous):

@saifoo.khan can you please help me

OpenStudy (saifoo.khan):

8x + 15 = 10x Bring the "x"s on one side of the equation. 8x - 10x + 15 = 0 Now take 15 on the other side, 8x - 10x = -15 Simplify..